In the year 2025, NASA was able to find another planet (XXV) which may be able to sustain




life. In order to determine if the planet’s water can sustain an earth organism’s life, Winkler




method was employed to determine dissolved oxygen content. A water sample from planet




XXV was obtained and then treated with the usual reagents for Winkler titration. A 50.0 mL




aliquot of this sample was titrated with the 0.010196M Na2S2O3 standard solution. It




required 4.16 mL of the titrant to reach the starch endpoint.




A. Calculate the amount of dissolved oxygen in the sample in terms of ppm O2. (FW O2 = 32.00)

1
Expert's answer
2022-05-25T18:44:03-0400

When you look at the reactions taking place in the Winkler method, it boils down to the fact that ONE MOLE of dissolved oxygen (O2) reacts with TWO MOLES of Na2S2O3 to produce the starch endpoint. Using this information, we can find the dissolved oxygen as follows ...


mols Na2S2O3 used = 4.16 mls x 1 L / 1000 mls x 0.010196 mols / L = 4.242x10-5 moles

mols O2 present in 50.0 mls = 4.242x10-5 mols Na2S2O3 x 1 mol O2 / 2 mols Na2S2O3 = 2.121x10-5 mols

mg O2 in 50.0 mls = 2.121x10-5 mols O2 x 32.0 g / mol x 1000 mg / g = 0.6787 mg O2 / 50.0 mls

Converting this to ppm O2 (which is equivalent to mg / L), we have ...

0.6787 mg O2 / 50.0 mls x 1000 mls / L = 13.6 mg / L = 13.6 ppm (3 sig. figs.)